im trying to figure out what to do with the remander...
the problem is
15 5
- -
10 H
i know i must cross multiply and i get 15H and 50 then divide but i get a remander of 5.. i dont know what to do with the remainder

Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)

If the problem is
15/10=5/h
then you get
15h=50
h=50/15
h=10/3=3 1/3

Answer by solver91311(24713)



Turn it into a fraction where the remainder is the numerator and the divisor is the denominator. Then reduce to lowest terms.




John